# Dispersion in Analogue Gravity

### v1

@@ Abstract
Conversely, a spacetime geometry with metric (1) can be viewed as a moving medium. This is true for arbitrary spacetime geometries, not only for Gordon’s metric (1): they all act on light like effective — possibly nonisotropic — moving media [6], and so does gravity, much like Newton initially thought.
@@
Unruh’s analogy gained prominence with the resolution of the trans-Planckian problem. In the standard Hawking derivation [8] modes near the horizon get exponentially blueshifted, implying arbitrarily large frequencies beyond the scale of our current theory. Hawking radiation appears to have come from regimes where we cannot trust physics anymore. Jacobson and Unruh showed [9, 10, 11] that realistic dispersive effects at high frequencies, far from invalidating the effect, actually preserve the Hawking-like radiation. Replacing the linear dispersion relation $\omega=c\,k$ by something like
@@
prevents unbounded frequency growth. Numerical [10] and analytic work [9, 11] indicated that once a horizon structure is present, the emission of Hawking radiation remains robust. The field of analogue gravity was born, had shown its promise, but was still an obscure subject at the fringe of the mainstream. This changed in the era of experiments.
@@
Throughout history, each significant leap in our understanding of gravity (from Newton’s mechanical viewpoint to Einstein’s geometric perspective) has been guided by powerful analogies. It now appears likely that future breakthroughs will also draw upon insights from analogue models, which points to deeper underlying structures beyond the apparent spacetime continuum, raising the question of how far these analogies can be extended and whether fully self–consistent Einstein equations can indeed be encoded in the language of a medium. Physics endeavors to present a unified portrayal of reality, but our understanding is inevitably entangled with the analogies that inspired it. And, in the end, who can definitively say where an analogue system ceases to be a mere analogy and becomes an essential guide to the deeper layers of gravitation?

### v2

@@ Abstract
Conversely, a spacetime geometry with metric (1) can be viewed as a moving medium. This is true for arbitrary spacetime geometries, not only for Gordon’s metric (1): they all act on light like effective — possibly nonisotropic — moving media [6], and so does gravity, much like Newton initially thought. [^1] This discussion applies to bosonic analogues (e.g., phonons/photons), while fermionic analogues require a tetrad structure [7].
@@
Unruh’s analogy gained prominence with the resolution of the trans-Planckian problem. In the standard Hawking derivation [9] modes near the horizon get exponentially blueshifted, implying arbitrarily large frequencies beyond the scale of our current theory. Hawking radiation appears to have come from regimes where we cannot trust physics anymore. Jacobson and Unruh showed [10, 11, 12] that realistic dispersive effects at high frequencies, far from invalidating the effect, actually preserve the Hawking-like radiation. Replacing the linear dispersion relation $\omega=c\,k$ by a higher-order one like
@@
which, though not universal [13], prevents unbounded frequency growth. Numerical [11] and analytic work [10, 12] indicated that once a horizon structure is present, the emission of Hawking radiation remains robust. The field of analogue gravity was born, had shown its promise, but was still an obscure subject at the fringe of the mainstream. This changed in the era of experiments.
@@
Throughout history, each significant leap in our understanding of gravity (from Newton’s mechanical viewpoint to Einstein’s geometric perspective) has been guided by powerful analogies. It now appears likely that future breakthroughs will also draw upon insights from analogue models, which points to deeper underlying structures beyond the apparent spacetime continuum, raising the question of how far these analogies can be extended and whether fully self–consistent Einstein equations can indeed be encoded in the language of a medium [36] . Physics endeavors to present a unified portrayal of reality, but our understanding is inevitably entangled with the analogies that inspired it. And, in the end, who can definitively say where an analogue system ceases to be a mere analogy and becomes an essential guide to the deeper layers of gravitation?
Acknowledgements
We thank Uwe Fischer and Grigory Volovik for helpful comments. E.E.E. gratefully acknowledges support of the the Weizmann Institute of Science. U.L. is supported by the Murray B. Koffler Professorial Chair.